OpenStudy (anonymous):

Derivative e^-2x*ln(1-2x)

OpenStudy (tkhunny):

Is there an actual problem statement? What have you done? Have you done one sort of like this, maybe from the same section?

OpenStudy (anonymous):

the derivative of the aforementioned expression

OpenStudy (anonymous):

http://gyazo.com/4beda3c73e097e7696e128375c7581ce

OpenStudy (tkhunny):

Okay, now the next two questions. It's a produce rule and a chain rule.

OpenStudy (tkhunny):

Can you find \(\dfrac{d}{dx}e^{-2x}\)? Can you find \(\dfrac{d}{dx}\ln(1-2x)\)?